# FIGRET: Fine-Grained Robustness-Enhanced Traffic Engineering

### v1

@@ 5.3. Coping with network failures
Figure 5 shows the performance of FIGRET, when different numbers of randomly selected links fail within the GEANT topology. We conduct a comparison between FIGRET, DOTE, Desensitization-based TE (Des TE), and Des TE that is also fault-aware, knowing the links that will fail in the future (FA Des TE). Their result is normalized against an oracle that possesses the best knowledge of both future demands and failures. Our results demonstrate that FIGRET achieves high resiliency to link failures. It outperforms both the DOTE and Des TE and is equally competitive with the FA Des TE, which has Oracle access to future failures. Our interpretation for this is that first, compared to DOTE, FIGRET, by incorporating robustness constraints, ensures that the traffic load on a path does not become excessive, thereby mitigating the impact of link failures on performance. The reason why FIGRET can compete with the performance of FA Des TE is due to the more accurate $D^{\text{expect}}$ obtained through deep learning and the benefits brought by fine-grained robustness enhancement, which offset the confusion induced by link failures. Our results on other topologies are presented in Appendix E. These results also exhibit similar characteristics.
@@
5.4. Robustness to demand changes
Temporary changes in traffic. We test FIGRET for traffic variability. Our method for increasing variability is as follows: For each source-destination (SD) pair in the real traffic, we generate a traffic fluctuation using the Gaussian distribution $N(\mu,\sigma^{2})$ multiplied by a factor $\alpha$, where $\mu=0$ and $\sigma$ is the standard deviation of the traffic demand $d_{st}$. Here, $\alpha$ is chosen from the set {0.2,0.5,1.0,2.0}, indicating the amplitude of the fluctuation. We observed how much the performance of FIGRET decreased after imposing sudden traffic changes, compared to its performance without such changes. The results are summarized in Table 3. It can be observed that when the factor $\alpha$ is relatively small, there is no significant decrease in FIGRET’s performance. At the same time, when the factor $\alpha=2$, introducing noise twice that of $N(0,\sigma^{2}_{D_{sd},[1-T]})$, the performance decrease does not exceed $20\%$.
@@ 5.4. Robustness to demand changes
Natural drift in traffic. We test the impact of natural traffic shifts on FIGRET. In the aforementioned experiment, we sorted the data chronologically, using the first 75% for training and the latter 25% for testing. However, in this section, we conduct training separately with 0%-25%, 25%-50%, and 50%-75% of the data, followed by testing on the last 25%. This approach allows us to observe the performance decline compared to when 75% of the data is used for training. We focus on two aspects: firstly, the impact of training with less data on FIGRET’s performance; secondly, the effect of the update frequency on performance. The results are summarized in Table 4. It is observed that FIGRET’s performance remains largely unaffected even a long time after training completion (exceeding three times the total duration of the training data). To fully understand how traffic data evolves across time, we also visually analyzed the data distribution across various times in the topology, presented in Appendix F. The analysis indicates that neither PoD-level nor ToR-level data exhibit drastic changes in their traffic patterns over time. However, the shift effect at the ToR level is somewhat more pronounced than at the PoD level. This corresponds with our results in Table 4. Through this analysis, we are inspired to recognize that the training of FIGRET does not necessarily need to be especially frequent.
@@
Table 4. Performance decline with natural drift in traffic: negative values indicate no degradation.
@@ 5.5. Interpreting FIGRET’s effectiveness
(a) Desensitization-based TE, PoD-level, Facebook DB cluster.
(b) FIGRET, PoD-level, Facebook DB cluster.
(c) Desensitization-based TE, ToR-level, Facebook DB cluster.
(d) FIGRET, ToR-level, Facebook DB cluster.
Figure 6. Compare the distribution of path sensitivities in Desensitization-based TE with FIGRET. Each point in the graph represents a path, with the x-axis indicating the variance of the traffic served by that path. The y-axis represents the average sensitivity of the path across all tested traffic demands.
Adaptive sensitivity analysis in FIGRET. We evaluate FIGRET’s ability to ascertain reasonable sensitivity constraints, as denoted by $\mathcal{F}((s,t))$ in §4. We compare the network configurations of Google’s Desensitization-based TE and FIGRET, examining the relationship between path sensitivity and traffic characteristics. The results are presented in Figure 6. When creating Figure 6, we normalize the variance of the traffic on the x-axis. For the y-axis, in calculating path sensitivity, we normalize the capacity of the edges (considering the edge with the smallest capacity in the topology as 1). Furthermore, for the ToR level topology in Figure 6(c) and 6(d), with 155 nodes representing $155\times 154$ SD pairs and three paths per SD pair, there are a total of 1551543 paths. Due to the excessive number of points to plot, we uniformly select $10\%$ of the SD pairs. From these results, we have the following observations: 1) For Desensitization-based TE, the path sensitivity of all paths, whether traffic is bursty or not, is constrained within a single path sensitivity constraint. 2) For FIGRET, path sensitivity settings depend on traffic characteristics: it’s dispersed for non-bursty and concentrated for bursty traffic. This enables detailed robustness analysis, achieving a better balance between performance and robustness to bursty traffic.
@@
Is deep learning essential for implementing the concept of fine-grained robustness enhancement? In the design of FIGRET, we harness deep learning to devise a robustness enhancement scheme based on complex traffic characteristics and network topologies. However, considering that commercial network operators may be reluctant to overhaul their existing in-production TE systems completely, we discuss in this section whether slight modifications to the existing TE schemes, incorporating the principles of robustness enhancement, could result in better performance compared to the original TE schemes. To investigate this, we make slight modifications to the TE system currently in use by Google’s Jupiter. We discover that by designing simple heuristic-based $\mathcal{F}$ path sensitivity constraint functions that align with fine-grained robustness concepts, the performance of TE algorithms can be improved. We employ linear functions that impose stricter constraints as traffic variability increases, and piecewise functions that divide traffic into more stable and bursty segments, applying looser and stricter constraints respectively. We validate that even such simple functions can improve the original TE’s performance to some extent (see Appendix C for more details).
Although it is feasible to directly apply fine-grained robustness enhancements to original systems, this approach encounters two significant issues: firstly, the use of LP methods for solving is notably slow. Secondly, manually configured functions often fall short of achieving optimal performance due to their inability to fully adapt to complex networks and dynamic traffic demands. Given these considerations, this paper employs a method based on deep learning.
@@
Traffic engineering. Traffic engineering has been extensively studied in various scenarios, including wide area networks (WANs) (Jain et al., 2013; Jiang et al., 2009; Kandula et al., 2005; Kandula et al., 2014; Zhang et al., 2015; Kumar et al., 2018b; Elwalid et al., 2001; Fortz and Thorup, 2000; Zhong et al., 2021; Wang et al., 2006; Zhang and Ge, 2005; Roughan et al., 2003; Applegate and Cohen, 2003; Suchara et al., 2011; Applegate et al., 2004; Bernárdez et al., 2021; Xu et al., 2023) and data center networks (Cao et al., 2023; Poutievski et al., 2022; Alizadeh et al., 2014; Benson et al., 2011; Al-Fares et al., 2010; Teh et al., 2020; Cao et al., 2021; Chen et al., 2018). Recently, SDN-based Traffic engineering has attracted considerable attention (Jain et al., 2013; Poutievski et al., 2022; Akyildiz et al., 2014; Agarwal et al., 2013; Zaicu et al., 2021; Bogle et al., 2019; Hong et al., 2013; Kumar et al., 2015; Liu et al., 2014).
Robustness-enhanced-based TE. A major focus of research on TE is to improve the algorithm’s robustness in handling sudden traffic bursts. Oblivious TE (Applegate and Cohen, 2003) aims to enhance robustness by consistently considering the most congested scenario, optimizing for the worst-case MLU among all possible DMs. COPE (Wang et al., 2006) optimizes MLU across a set of DMs predicted by historical DMs while holding a worst-case MLU guarantee. Desensitization-based TE (Poutievski et al., 2022; Teh et al., 2020), by introducing the concept of path sensitivity which characterizes the impact of traffic bursts on the path, aims to enhance robustness by optimizing the MLU while ensuring low path sensitivity across a historical set of DMs. Robustness-enhanced TE schemes typically exhibit coarse granularity by treating each source-destination pair in the traffic matrix equally. In contrast, FIGRET stands apart by employing fine-grained robustness enhancement, resulting in enhanced performance.
Machine learning-based TE. Machine learning-based TE can be broadly categorized into two classes. The first class is Demand Prediction-based TE (Namyar et al., 2022; Abuzaid et al., 2021; Kumar et al., 2018a; Narayanan et al., 2021), where these methods utilize predictive models to initially compute an explicitly predicted next Traffic Matrix (TM). Subsequently, optimization is performed based on the predicted TM. The second class replaces explicit demand prediction with end-to-end optimization, directly mapping a recent historical window of TMs to TE configurations (Perry et al., 2023; Valadarsky et al., 2017). Despite harnessing the powerful capabilities of deep learning to learn relationships between traffic flows, both of these classes overlook the consideration of unexpected traffic bursts.
@@
In this work, we introduce FIGRET, a new design point for TE that enhances robustness in a fine-grained manner based on traffic characteristics. FIGRET employs path sensitivity to manage traffic bursts and utilizes a deep learning strategy with a well-designed loss function to produce TE solutions that consider robustness at a fine-grained level. Our experimental results demonstrate that FIGRET achieves an effective balance between average performance and robustness, resulting in high-quality TE solutions.
@@ C.1. Linear function
We apply the parameters listed in Table 6 and conduct tests on the PoD-level Facebook DB. The results are summarized in Figure 8. As shown in Figure 8, the application of Strategy 1, which involves imposing stricter path sensitivity constraints (as in the case of the group {1,2,3}), leads to an enhanced capability to handle bursty traffic. On the other hand, with the implementation of Strategy 2, an improvement in average performance is observed, as demonstrated by the comparison of groups {3,4}.
@@ C.2. Piecewise function
In our study, we conducted a series of tests on the parameters listed in Table 7 on PoD-level Facebook DB, with the results summarized in Figure 10. As illustrated in the figure, with fixed values of Min and Max, a larger Breakpoint setting enhances the average performance of TE methods. This trend is evident in groups {1,2,3} and {5,6,7}, where an increase in Breakpoint corresponds to a lower average performance boxplot. Conversely, with a constant Breakpoint, maintaining Max while reducing Min enhances the TE methods’ ability to manage bursty traffic, as shown in the comparison of group {1,4}. Similarly, fixing Min while increasing Max leads to better average performance of TE methods, as observed in the comparison between groups {4,5}.
@@ D.2. The inappropriateness of CNN
In Figure 11, there is a source $s$ and three destinations $t_{1},t_{2},t_{3}$, with a convolution kernel size of 2. This kernel aims to extract local information of length 2 from the traffic demands. However, extracting local information between ${d_{1},d_{2}}$ and ${d_{2},d_{3}}$ proves ineffective because $p_{st_{1}}$ and $p_{st_{2}}$, as well as $p_{st_{2}}$ and $p_{st_{3}}$, do not share a common edge. Conversely, $p_{st_{1}}$ and $p_{st_{3}}$ do share a common edge, and their traffic demands should be considered jointly. This limitation cannot be addressed by such convolution layers that are designed to extract local information. Unlike image data, which CNNs excel at handling due to the localized relevance of information, TE problems do not exhibit a distinct need for extracting local information because traffic demands within a convolutional locality may not necessarily pass through a common edge, making the use of convolution for local information extraction minimally effective. Therefore, in our model design, we chose not to incorporate convolutional layers. Instead, we opted for directly utilizing fully connected layers.
(a) Network topology
(b) Diagram illustrating convolutional operation
Figure 11. Diagram illustrating the convolution operation in TE. Traffic demands for paths that share a common edge need to be considered jointly due to their potential impact on each other. In contrast, for paths without a shared edge, jointly considering their traffic demands is ineffective.

### v3

@@ 5.3. Coping with network failures
Figure 14 shows the performance of FIGRET, when different numbers of randomly selected links fail within the GEANT topology. We conduct a comparison between FIGRET, DOTE, Desensitization-based TE (Des TE), and Des TE that is also fault-aware (FA Des TE). Fault-aware refers to knowing the links that will fail in the future and optimizing only over the paths without failures, without the need to reroute the traffic around failed paths after the TE solutions are determined. Their result is normalized against an oracle that possesses the best knowledge of both future demands and failures. Our results demonstrate that FIGRET achieves high resiliency to link failures. It outperforms both the DOTE and Des TE and is equally competitive with the FA Des TE, which has Oracle access to future failures. Our interpretation for this is that first, compared to DOTE, FIGRET, by incorporating robustness constraints, ensures that the traffic load on a path does not become excessive, thereby mitigating the impact of link failures on performance. The reason why FIGRET achieves comparable performance with FA Des TE is twofold: firstly, Des TE’s explicit prediction of the traffic matrix introduces errors; secondly, it lacks fine-grained robustness enhancement. These factors offset the advantage FA Des TE gains from knowing link failures. Our results on other topologies are presented in Appendix E. These additional results also exhibit similar characteristics.
@@
Table 4. Performance decline with natural drift in traffic. Negative values indicate no degradation.
Network Factor
0.2 0.5 1.0 2.0
PoD DB average -0.3% 0.7% 4.8% 15.3%
90th Pct. 1.7% 1.0% 6.6% 27.9%
pFabric average 0.2% -0.6% -1.7% -2.7%
90th Pct. 0.5% 0.7% -0.4% 1.6%
ToR DB average 0.9% 2.7% 17.0% 38.6%
90th Pct. 0.0% 2.6% 23.1% 32.4%
Table 5. Performance decline under worst-case conditions. Negative values indicate no degradation.
5.4. Robustness to demand changes
Temporal changes in traffic. We test FIGRET for traffic variability. Our method for increasing variability is as follows: For each source-destination (SD) pair in the real traffic, we generate a traffic fluctuation using the Gaussian distribution $N(\mu,\sigma^{2})$ multiplied by a factor $\alpha$, where $\mu=0$ and $\sigma$ is the standard deviation of the traffic demand $d_{st}$. Here, $\alpha$ is chosen from the set {0.2,0.5,1.0,2.0}, indicating the amplitude of the fluctuation. We observe how much the performance of FIGRET decreased after imposing sudden traffic changes, compared to its performance without such changes. The results are summarized in Table 5. It can be observed that when the factor $\alpha$ is relatively small, there is no significant decrease in FIGRET’s performance. At the same time, when the factor $\alpha$ is 2, doubling the noise of $N(0,\sigma^{2}_{D_{sd},[1-T]})$, the performance does not decrease by more than $20\%$.
Natural drift in traffic. We test the impact of natural traffic shifts on FIGRET. In the aforementioned experiment, we sorted the data chronologically, using the first 75% for training and the latter 25% for testing. However, in this section, we conduct training separately with 0%-25%, 25%-50%, and 50%-75% of the data, followed by testing on the last 25%. This approach allows us to observe the performance decline compared to when 75% of the data is used for training. We focus on two aspects: firstly, the impact of training with less data on FIGRET’s performance; secondly, the effect of the update frequency on performance. The results are summarized in Table 5. It is observed that FIGRET’s performance remains largely unaffected even a long time after training completion (exceeding two times the total duration of the training data). To fully understand how traffic data evolves across time, we also visually analyzed the data distribution across various times, presented in Appendix F. The analysis indicates that neither PoD-level nor ToR-level data exhibit drastic changes in their traffic patterns over time. However, the shift effect at the ToR level is somewhat more pronounced than at the PoD level. This corresponds with our results in Table 5. Through this analysis, we are inspired that the FIGRET training does not necessarily need to be especially frequent.
Worst-case performance. We evaluate the performance of FIGRET under worst-case conditions. Considering that FIGRET develops strategies based on the characteristics of historical data, we intentionally reverse the order of the magnitude of temporal traffic fluctuations among SD pairs, applying larger traffic fluctuations to those pairs with historically lower variances. The results are summarized in Table 5. Compared to the results in Table 5, although there is indeed a greater decline in FIGRET’s performance, with PoD DB and ToR DB experiencing a 30%-40% reduction when $\alpha=2$, the performance did not completely fail. Additionally, we employ the Spearman rank correlation coefficient to assess the consistency of variance ranking changes between the test and train sets across different SD pairs within our dataset. The analysis reveals a high Spearman rank correlation coefficient, with PoD DB at 0.92 and ToR DB at 0.98. These results suggest that the actual occurrence of the worst-case scenario we are considering is rare. Furthermore, for the pFabric system, since the sources and destinations of traffic are selected through a uniform random process, the differences in variance among SD pairs are not pronounced. Therefore, the performance variations are not significant.
@@ 5.5. Interpreting FIGRET’s effectiveness
Sensitivity analysis in FIGRET. We numerically examine how FIGRET achieves fine-grained robustness. We compare the network configurations of Google’s hedge-based TE and FIGRET, examining the relationship between path sensitivity and traffic characteristics. The results are presented in Figure LABEL:fig:_interpret_termol. When creating Figure LABEL:fig:_interpret_termol, we normalize the variance of the traffic on the x-axis. For the y-axis, when calculating path sensitivity, we normalize the capacity of the edges (considering the edge with the smallest capacity in the topology as 1). Furthermore, for the ToR level topology in Figure LABEL:fig:jupiter_facebook_tor_a and LABEL:fig:figret_facebook_tor_a, with 155 nodes representing $155\times 154$ SD pairs and three paths per SD pair, there are a total of 1551543 paths. Due to the excessive number of points to plot, we uniformly select $10\%$ of the SD pairs for better visualization. From these results, we have the following observations: 1) For hedge-based TE, the path sensitivity of all paths, whether traffic is bursty or not, is constrained within a constant upper bound. 2) For FIGRET, path sensitivity settings depend on traffic characteristics: it’s dispersed for non-bursty and concentrated for bursty traffic. This enables fine-grained robustness analysis, achieving a better balance between normal-case performance and robustness to burst traffic.
@@
Is deep learning essential for implementing the concept of fine-grained robustness enhancement? In the design of FIGRET, we harness deep learning to devise a robustness enhancement scheme based on complex traffic characteristics and network topologies. However, considering that commercial network operators may be reluctant to overhaul their existing in-production TE systems completely, we discuss in this section whether slight modifications to the existing TE schemes, incorporating the principles of robustness enhancement, could result in better performance compared to the original TE schemes. To investigate this, we make slight modifications to the TE system currently in use by Google’s Jupiter data centers. We discover that by designing simple heuristic-based $\mathcal{F}$ path sensitivity constraint functions that align with fine-grained robustness concepts, the performance of TE algorithms can be improved. We use linear functions to impose stricter constraints as traffic variability increases. We also use piecewise functions to divide traffic into more stable and bursty segments, applying looser constraints to stable segments and stricter constraints to bursty segments. We validate that even such simple functions can improve the original TE’s performance. For example, when using linear functions, if we maintain the constraints for high bursty traffic but relax the constraints for low bursty traffic, we can achieve a 5% reduction in the average normal-case MLU while maintaining robustness (see Appendix C for more details).
Although it is feasible to directly apply fine-grained robustness enhancements to traditional TE systems currently in operation, this approach encounters two significant issues: firstly, the use of LP methods for solving is notably slow. Secondly, manually configured functions often fall short of achieving optimal performance due to their inability to fully adapt to complex networks and dynamic traffic demands. Given these considerations, this paper employs a method based on deep learning.
@@
Is it feasible to deploy deep learning-based TE methods in practice? In general, there are two main obstacles that may prevent the practical deployment of deep learning-based methods: 1) the learning-based methods may sometimes produce incorrect or infeasible solutions, and 2) the performance results are not easily interpretable, raising concerns about the solution reliability in different cases. Nevertheless, we argue that the above issues are less of a concern for FIGRET. First, the constraints of TE, which require that the sum of proportions served by all paths for a given source-destination pair equals one, can be easily enforced by normalizing the outputs of the neural network, thus avoiding infeasible solutions. Second, FIGRET’s superior performance is explainable, as demonstrated in §5.5.
Can the concept of fine-grained robustness be extended to other objectives? Firstly, fine-grained robustness in traffic not only applies to minimizing MLU but can also be extended to other objectives such as minimizing latency. For instance, selecting the shortest path for non-burst traffic can achieve minimal latency; however, for potential burst traffic, employing multipath transmission may be necessary to avoid congestion on the shortest paths. Secondly, the concept of fine-grained robustness is also applicable to handling link failures. Existing strategies, such as FFC (Liu et al., 2014), often optimize for all possible failure scenarios, which can lead to low utilization and over-provisioning. By adopting a fine-grained approach that considers the probability of failure scenarios, resource utilization, and fault management can be balanced more effectively.
@@
Robustness-enhanced-based TE. A major focus of research on TE is to improve the algorithm’s robustness in handling sudden traffic bursts. Oblivious TE (Applegate and Cohen, 2003) aims to enhance robustness by consistently considering the most congested scenario, optimizing for the worst-case MLU among all possible DMs. COPE (Wang et al., 2006) optimizes MLU across a set of DMs predicted by historical DMs while holding a worst-case MLU guarantee. Desensitization-based TE (Poutievski et al., 2022; Teh et al., 2020), by introducing the concept of path sensitivity which characterizes the impact of traffic bursts on the path, aims to enhance robustness by optimizing the MLU while ensuring low path sensitivity. Robustness-enhanced TE schemes typically exhibit coarse granularity by treating each source-destination pair in the traffic matrix equally. In contrast, FIGRET stands apart by employing fine-grained robustness enhancement, resulting in a better balance between robustness and average-case performance.
Threshold-based VLB Routing: TROD (Cao et al., 2021, 2023) introduces a threshold-based routing strategy in optical data centers to handle unexpected traffic bursts. The basic idea is that when traffic demand is below a predetermined threshold, the traffic is routed via shortest paths. When traffic demand exceeds the threshold, the burst traffic is load-balanced across all non-shortest paths. TROD maintains normal-case performance by routing most traffic via the shortest paths, while it enhances robustness by dispersing bursts across multiple paths. Hence, TROD strikes a good balance between normal-case and worst-case performance. Unfortunately, TROD’s routing algorithm requires special hardware support. P4 switches can support TROD’s routing algorithm, but they are not widely deployed in data centers. In contrast, FIGRET does not require specialized hardware and only needs switches that support WCMP.
Machine learning-based TE. Machine learning-based TE can be broadly categorized into two classes. The first class is Demand Prediction-based TE [1, 31, 35, 36]. These methods use predictive models to estimate the next traffic matrix (TM) and then perform optimization based on the estimated TM. The second class replaces explicit demand prediction with end-to-end optimization, directly mapping a recent historical window of TMs to TE configurations (Perry et al., 2023; Valadarsky et al., 2017). Despite harnessing the powerful capabilities of deep learning to learn relationships between traffic flows, both of these classes overlook the consideration of unexpected traffic bursts.
Network planning. The objective of network planning is to configure network capacity to ensure good performance for all possible traffic patterns. Consequently, Meta models the traffic based on the Hose model (Eason et al., 2023; Ahuja et al., 2021) and optimizes for the worst-case performance. This approach is not directly applicable to TE because it adjusts network capacity rather than routing weights on paths. Furthermore, even if one may repurpose a network planning approach to solve TE problems, the resulting solution lacks the capability to balance normal-case performance and burst-case performance.
@@
In this work, we introduce FIGRET, a new design point for TE that enhances robustness in a fine-grained manner based on traffic characteristics. FIGRET employs path sensitivity to manage traffic bursts and utilizes a deep learning strategy with a well-designed loss function to produce TE solutions that consider robustness at a fine-grained level. Our experimental results demonstrate that FIGRET achieves an effective balance between normal-case performance and robustness, resulting in high-quality TE solutions.
ACKNOWLEDGMENTS
We sincerely thank our shepherd Manya Ghobadi and the anonymous reviewers for their constructive feedback. This work was supported by the NSF China (No. 62272292, No. 62132009 and No. 61960206002).
@@ ACKNOWLEDGMENTS
Appendices are supporting material that has not been peer-reviewed.
@@ C.1. Linear function
We apply the parameters listed in Table 7 and conduct tests on the PoD-level Meta DB. The results are summarized in Figure 19. As shown in Figure 19, the application of Strategy 1, which involves imposing stricter path sensitivity constraints (as in the case of the group {1,2,3}), leads to an enhanced capability to handle bursty traffic. On the other hand, with the implementation of Strategy 2, an improvement in average performance is observed, as demonstrated by the comparison of groups {3,4}. Comparing parameters 5 with parameters 1, we observe that by maintaining the sensitivity constraint for high bursty traffic while relaxing the sensitivity constraint for non-bursty traffic, parameters 5 can reduce the normal-case performance (below the 75th percentile) MLU by 5% at the expense of increasing the peak MLU by 0.7%.
@@ C.2. Piecewise function
In our study, we conducted a series of tests on the parameters listed in Table 8 on PoD-level Meta DB, with the results summarized in Figure 21. As illustrated in the figure, with fixed values of Min and Max, a larger Breakpoint setting enhances the average performance of TE methods. This trend is evident in groups {1,2,3} and {5,6,7}, where an increase in Breakpoint corresponds to a lower average performance boxplot. Conversely, with a constant Breakpoint, maintaining Max while reducing Min enhances the TE methods’ ability to manage bursty traffic, as shown in the comparison of group {1,4}. Similarly, fixing Min while increasing Max leads to better average performance of TE methods, as observed in the comparison between groups {4,5}.
@@ D.2. The inappropriateness of CNN
In Figure LABEL:fig:_cnn, there is a source $s$ and three destinations $t_{1},t_{2},t_{3}$, with a convolution kernel size of 2. This kernel aims to extract local information of length 2 from the traffic demands. However, extracting local information between ${d_{1},d_{2}}$ and ${d_{2},d_{3}}$ proves ineffective because $p_{st_{1}}$ and $p_{st_{2}}$, as well as $p_{st_{2}}$ and $p_{st_{3}}$, do not share a common edge. Conversely, $p_{st_{1}}$ and $p_{st_{3}}$ do share a common edge, and their traffic demands should be considered jointly. This limitation cannot be addressed by such convolution layers that are designed to extract local information. Unlike image data, which CNNs excel at handling due to the localized relevance of information, TE problems do not exhibit a distinct need for extracting local information because traffic demands within a convolutional locality may not necessarily pass through a common edge, making the use of convolution for local information extraction minimally effective. Therefore, in our model design, we chose not to incorporate convolutional layers. Instead, we opted for directly utilizing fully connected layers.